# A Posteriori Single- and Multi-Goal Error Control and Adaptivity for Partial Differential Equations

### v2

@@ Abstract
This chapter reviews goal-oriented a posteriori error control, adaptivity and solver control for finite element approximations to boundary and initial-boundary value problems for stationary and non-stationary partial differential equations, respectively. In particular, coupled field problems with different physics may require simultaneously the accurate evaluation of several quantities of interest, which is achieved with multi-goal oriented error control. Sensitivity measures are obtained by solving an adjoint problem. Error localization is achieved with the help of a partition-of-unity. We also review and extend theoretical results for efficiency and reliability by employing a saturation assumption. The resulting adaptive algorithms allow to balance discretization and non-linear iteration errors, and are demonstrated for four applications: Poisson’s problem, non-linear elliptic boundary value problems, stationary incompressible Navier-Stokes equations, and regularized parabolic $p$-Laplace initial-boundary value problems. Therein, different finite element discretizations in two different software libraries are utilized, which are partially accompanied with open-source implementations on GitHub.
@@ 1 Introduction
This chapter is devoted to a goal-oriented a posteriori error control of single and multiple quantities of interest (QoI) and Adaptive Finite Element Methods (AFEM) for solving stationary and non-stationary, linear and non-linear partial differential equations (PDEs) and systems of PDEs. AFEM were certainly inspired by the pioneer work [13] of Babuška and Rheinboldt. Further important studies on a posteriori error controlled adaptive finite element methods are [85, 65, 167, 43, 211] to name a few. The first comprehensive monograph [223] by Verfürth was another AFEM milestone. We refer to the survey papers [105], [35], [67], and to the monographs [4], [15], [18], [125], [170], [192], [106], [193] for an overview of a posteriori error estimates and adaptive finite element techniques.
The governing discretization in this chapter is the finite element method [54, 51, 71, 136, 62, 107], but isogeometric analysis could be used in a similar fashion [138, 73, 36]. We consider stationary settings as well as time-dependent cases, modeled as space-time problems and solved by space-time discretization methods; see, e.g., [140, 137, 220, 209, 172, 210, 148, 215, 149] and the references therein.
@@
Conceptional developments on goal-oriented error estimation include a safe-guarded method established in [174] and guaranteed bounds derived in [5]. Abstract analyses and reformulations were presented in [110, 142]. First convergence rates of goal-oriented error estimation were obtained in [166], weighted marking [28], bounding techniques for goal-oriented error estimates for elliptic problems [147], goal oriented flux reconstruction [168], goal-oriented error estimation for the fractional-step-theta scheme [158], a partition-of-unity localization including effectivity estimates for single goal functionals [198], partition-of-unity localization for multiple goal functionals [103], space-time partition-of-unity localization [96, 219, 102], equilibrated flux [52], linearization errors [122]. Theoretical work showing optimality of adaptive algorithms was summarized in four axioms of adaptivity in [63] and dates back to residual-based a posteriori error estimates in [85]. First rigorous convergence results for goal-oriented estimators go back to [166, 28], then [110, 135, 134], and recent findings are [31, 27, 29]. We also mention our own prior work on such theoretical advancements that include effective estimates with common upper bounds of the dual-weighted residual estimator and the error indicators [198]. Efficiency and reliability estimates for the dual-weighted residual method using a saturation assumption were shown in [100], and smart algorithms switching between solving high-order adjoint problems or using interpolations were derived in [101].
@@
Single- and multigoal-oriented error control and adaptivity has found numerous applications. These include Poisson’s problem [34, 33, 180, 103, 61], stationary linear elasticity [185], nonlinear finite deformation elasticity [121], space-time adaptivity for parabolic equations [204], space-time elasticity and the wave equation [182, 16], multi-rate discretizations of coupled parabolic/hyperbolic problems [208], $p$-Laplace problems [99, 101, 96, 190], elasto-plasticity [186, 187, 188] and visco-plasticity [155], incompressible flow [30, 50, 37, 38, 97, 100, 101, 201], transport problems [146], transport problems with coupled flow [22, 57, 58], Boussinesq equations and coupled flow with temperature [41, 40, 93], reactive flows [26, 194, 49], biomechanics [90, 60], uncertain inputs [39], Maxwell’s equations [139], elliptic eigenvalue problems [129, 72], modeling errors [175, 48, 183], applications to polygonal meshes [226], conforming and nonconforming approximations with inexact solvers [154], anisotropic mesh refinement [195], heterogeneous multiscale problems [70, 152, 153, 150, 84], fluid-structure interaction [123, 196, 87, 88, 89, 227, 228, 112, 222, 197, 109, 3], phase-field slits and fracture [229, 230, 231], sea ice simulations [155], optimal control [32, 25, 160, 156, 132, 224, 133, 189, 131, 95], obstacle/contact problems and variational inequalities [45, 46, 44, 205, 184, 214, 230], finite cell methods [213, 81], balancing discretization and iteration errors [157, 191, 190, 99, 82, 83], financial mathematics [118], goal-oriented model order reduction [162, 113, 114, 141, 66], neural network enhanced dual-weighted residual technologies [56, 165, 200], adaptive multiscale predictive modeling [176], and open-source software developments for goal-oriented error estimation [119] [^5] https://github.com/winnifried/dopelib, [61] [^6] https://github.com/jhale/fenicsx-error-estimation, [90] [^7] https://doi.org/10.6084/m9.figshare.8128178.v1, [145] [^8] https://github.com/dtm-project/dwr-diffusion, [219, 216] [^9] https://github.com/jpthiele/pu-dwr-diffusion, [^10] https://github.com/jpthiele/pu-dwr-combustion.
@@
The outline of this book chapter is as follows. In Section 2, the notation and abstract setting are introduced. Then, in Section 3, single goal-oriented error control is discussed. Section 4 focuses on error estimation for non-standard discretizations. Next, in Section 5, multigoal-oriented error estimators are explained. In Section 6, three applications are considered, and computationally analyzed. The book chapter is concluded in Section 7, and some current research directions are outlined.
@@ 2 Notation, Abstract Setting, Finite Element Discretization
Throughout this book chapter, let $d$ be the space dimension. We denote by $Q:=\Omega\times I\subset\mathbb{R}^{d+1}$ the space-time domain (cylinder), where $\Omega\subset\mathbb{R}^{d}$ is the bounded and Lipschitz spatial domain with the boundary $\Gamma=\partial\Omega$, and $I:=(0,T)$ denotes the temporal domain (time interval). Here, $T>0$ is the end time value. Furthermore, we use the standard notations for Lebesgue, Sobolev, and Bochner spaces like $L_{p}(\Omega)$, $W_{p}^{k}(\Omega),\mathring{W}_{p}^{k}(\Omega),H^{k}(\Omega)=W_{2}^{k}(\Omega),H^{k}_{0}(\Omega)=\mathring{W}_{2}^{k}(\Omega)$, $L_{p}(0,T;\mathring{W}_{p}^{1}(\Omega))$ etc.; see, e.g., [51, 54, 232].
@@ 2.2.2 Finite Elements on Hypercubes $Q_{k}$
To facilitate adaptive mesh refinement and to avoid connecting elements, we use the concept of hanging nodes. Elements are allowed to have nodes that lie on the midpoints of the faces or edges of neighboring cells. In our implementations, at most, one hanging node is allowed on each face or edge. In three dimensions, this concept is generalized to subplanes and faces because we must deal with two types of lower manifolds. To enforce global continuity (i.e., global conformity), the degrees of freedom located on the interface between different refinement levels have to satisfy additional constraints. They are determined by interpolation of neighboring degrees of freedom. Therefore, hanging nodes do not carry any degrees of freedom. For more details on this, we refer to [62]. To ensure the approximation properties of the finite element spaces, additional conditions on the geometry of the cells are required. The two classical assumptions from the literature, e.g., [54, 51] are the so-called uniformity and the weaker concept of quasi-uniformity.
@@ 3.2 Adjoint Problem
In order to realize goal-oriented error control in this book chapter, we use the DWR method, which requires solving an adjoint problem. This adjoint formulation follows from the Lagrange formalism and is given by: Find $z\in V$ such that
@@ 3.3 An Error Identity
$\rho^{*}(\tilde{u},\tilde{z})(\cdot)$ $:=J^{\prime}(u)-\mathcal{A}^{\prime}(\tilde{u})(\cdot,\tilde{z}),$ (7)
@@
$\mathcal{L}^{\prime}(\tilde{x})(x-\tilde{x})=$ $\underbrace{J^{\prime}(\tilde{u})(e)-\mathcal{A}^{\prime}(\tilde{u})(e,\tilde{z})}_{=\rho^{*}(\tilde{u},\tilde{z})(u-\tilde{u})}\underbrace{-\mathcal{A}(\tilde{u})(e^{*})}_{=\rho(\tilde{u})(z-\tilde{z})},$
@@ 3.4 Error Estimation in Enriched Spaces
$\begin{split}\mathcal{R}^{(3)(2)}:=\frac{1}{2}\int_{0}^{1}&[J^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)})-\mathcal{A}^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)},\tilde{z}+se_{h}^{*(2)})\\ &-3\mathcal{A}^{\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)})]s(s-1)\,ds,\end{split}$ (13)
@@
holds, where $\rho(\tilde{u})(\cdot):=-\mathcal{A}(\tilde{u})(\cdot)$ and $\rho^{*}(\tilde{u},\tilde{z})(\cdot):=J^{\prime}(u)-\mathcal{A}^{\prime}(\tilde{u})(\cdot,\tilde{z})$.
@@
where $\rho_{h}(\tilde{u})(\cdot):=-\mathcal{A}^{(2)}(\tilde{u})(\cdot)$, $\rho^{*}_{h}(\tilde{u},\tilde{z})(\cdot):=J^{\prime}_{h}(u)-(\mathcal{A}^{(2)})^{\prime}(\tilde{u})(\cdot,\tilde{z})$, and
$\begin{split}\mathcal{R}_{h}^{(3)}:=\frac{1}{2}\int_{0}^{1}&[J_{h}^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)})-(\mathcal{A}^{(2)})^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)},\tilde{z}+se_{h}^{*(2)})\\ &-3(\mathcal{A}^{(2)})^{\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)})]s(s-1)\,ds,\end{split}$
@@ 3.5.3 The Discretization Error Estimator $\eta_{h}^{(2)}$
$\frac{1}{1+b_{h}}\left(|\eta_{h}^{(2)}|-|\eta_{k}+\eta_{\mathcal{R}}^{(2)}|\right)\leq|J(u)-J(\tilde{u})|\leq\frac{1}{1+b_{h}}\left(|\eta_{h}^{(2)}|+|\eta_{k}+\eta_{\mathcal{R}}^{(2)}|\right)$
@@
$\leq\frac{1}{1-b_{h}}|\eta^{(2)}|\leq\frac{1}{1+b_{h}}\left(|\eta_{h}^{(2)}|+|\eta_{k}+\eta_{\mathcal{R}}^{(2)}|\right).$
@@
$(1-\alpha_{\eta_{h}^{(2)}})\frac{1}{1+b_{h}}|\eta_{h}^{(2)}|\leq|J(u)-J(\tilde{u})|\leq(1+\alpha_{\eta_{h}^{(2)}})\frac{1}{1+b_{h}}|\eta_{h}^{(2)}|,$
@@ 3.6 Effectivity Indices for Enriched Spaces
$\frac{1}{1-b_{h}}|\eta^{(2)}|\leq|J(u)-J(\tilde{u})|\leq\frac{1}{1+b_{h}}|\eta^{(2)}|.$
@@
$\frac{1}{1-b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|\leq 1\leq\frac{1}{1+b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|.$
@@
$(1-\alpha_{\eta_{h}^{(2)}})\frac{1}{1+b_{h}}|\eta_{h}^{(2)}|\leq|J(u)-J(\tilde{u})|\leq(1+\alpha_{\eta_{h}^{(2)}})\frac{1}{1+b_{h}}|\eta_{h}^{(2)}|,$
@@
$\frac{1+\alpha_{\eta_{h}^{(2)}}}{1-b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|\leq 1\leq\frac{1-\alpha_{\eta_{h}^{(2)}}}{1+b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|.$

### v3

@@ Abstract
This work reviews goal-oriented a posteriori error control, adaptivity and solver control for finite element approximations to boundary and initial-boundary value problems for stationary and non-stationary partial differential equations, respectively. In particular, coupled field problems with different physics may require simultaneously the accurate evaluation of several quantities of interest, which is achieved with multi-goal oriented error control. Sensitivity measures are obtained by solving an adjoint problem. Error localization is achieved with the help of a partition-of-unity. We also review and extend theoretical results for efficiency and reliability by employing a saturation assumption. The resulting adaptive algorithms allow to balance discretization and non-linear iteration errors, and are demonstrated for four applications: Poisson’s problem, non-linear elliptic boundary value problems, stationary incompressible Navier-Stokes equations, and regularized parabolic $p$-Laplace initial-boundary value problems. Therein, different finite element discretizations in two different software libraries are utilized, which are partially accompanied with open-source implementations on GitHub.
@@ 1 Introduction
This work is devoted to a goal-oriented a posteriori error control of single and multiple quantities of interest (QoI) and Adaptive Finite Element Methods (AFEM) for solving stationary and non-stationary, linear and non-linear partial differential equations (PDEs) and systems of PDEs. AFEM were certainly inspired by the pioneer work [12] of Babuška and Rheinboldt. Further important studies on a posteriori error controlled adaptive finite element methods are [80, 60, 158, 42, 202] to name a few. The first comprehensive monograph [213] by Verfürth was another AFEM milestone. We refer to the survey papers [98], [34], [62], and to the monographs [4], [14], [17], [116], [161], [183], [99], [184] for an overview of a posteriori error estimates and adaptive finite element techniques.
The governing discretization in this work is the finite element method [52, 50, 66, 127, 57, 100], but isogeometric analysis could be used in a similar fashion [129, 68, 35]. We consider stationary settings as well as time-dependent cases, modeled as space-time problems and solved by space-time discretization methods; see, e.g., [131, 128, 210, 200, 163, 201, 139, 205, 140] and the references therein.
@@
Conceptional developments on goal-oriented error estimation include a safe-guarded method established in [165] and guaranteed bounds derived in [5]. Abstract analyses and reformulations were presented in [103, 133]. First convergence rates of goal-oriented error estimation were obtained in [157], weighted marking [27], bounding techniques for goal-oriented error estimates for elliptic problems [138], goal oriented flux reconstruction [159], goal-oriented error estimation for the fractional-step-theta scheme [149], a partition-of-unity localization including effectivity estimates for single goal functionals [189], partition-of-unity localization for multiple goal functionals [96], space-time partition-of-unity localization [90, 209], equilibrated flux [51], linearization errors [114]. Theoretical work showing optimality of adaptive algorithms was summarized in four axioms of adaptivity in [58] and dates back to residual-based a posteriori error estimates in [80]. First rigorous convergence results for goal-oriented estimators go back to [157, 27], then [103, 126, 125], and recent findings are [30, 26, 28]. We also mention our own prior work on such theoretical advancements that include effective estimates with common upper bounds of the dual-weighted residual estimator and the error indicators [189]. Efficiency and reliability estimates for the dual-weighted residual method using a saturation assumption were shown in [94], and smart algorithms switching between solving high-order adjoint problems or using interpolations were derived in [95].
@@
Single- and multigoal-oriented error control and adaptivity has found numerous applications. These include Poisson’s problem [33, 32, 171, 96], stationary linear elasticity [176], space-time adaptivity for parabolic equations [195], space-time elasticity and the wave equation [173, 15], multi-rate discretizations of coupled parabolic/hyperbolic problems [199], $p$-Laplace problems [93, 95, 90, 181], elasto-plasticity [177, 178, 179] and visco-plasticity [146], incompressible flow [29, 49, 36, 37, 91, 94, 95, 192], transport proplems [137], transport problems with coupled flow [21, 54, 55], Boussinesq equations and coupled flow with temperature [40, 39, 87], reactive flows [25, 185, 48], uncertain inputs [38], Maxwell’s equations [130], elliptic eigenvalue problems [120, 67], modeling errors [166, 47, 174], applications to polygonal meshes [216], conforming and nonconforming approximations with inexact solvers [145], anisotropic mesh refinement [186], heterogeneous multiscale problems [65, 143, 144, 141, 79], fluid-structure interaction [115, 187, 82, 83, 84, 217, 218, 105, 212, 188, 102, 3], phase-field slits and fracture [219, 220, 221], sea ice simulations [146], optimal control [31, 24, 151, 147, 123, 214, 124, 180, 122, 89], obstacle/contact problems and variational inequalities [44, 45, 43, 196, 175, 204, 220], finite cell methods [203, 76], balancing discretization and iteration errors [148, 182, 181, 93, 77, 78], financial mathematics [111], goal-oriented model order reduction [153, 106, 107, 132, 61], neural network enhanced dual-weighted residual technologies [53, 156, 191], adaptive multiscale predictive modeling [167], and open-source software developments for goal-oriented error estimation [112] [^5] https://github.com/winnifried/dopelib, [136] [^6] https://github.com/dtm-project/dwr-diffusion, [209, 206] [^7] https://github.com/jpthiele/pu-dwr-diffusion, [^8] https://github.com/jpthiele/pu-dwr-combustion.
@@
The outline of this work is as follows. In Section 2, the notation and abstract setting are introduced. Then, in Section 3, single goal-oriented error control is discussed. Section 4 focuses on error estimation for non-standard discretizations. Next, in Section 5, multigoal-oriented error estimators are explained. In Section 6, three applications are considered, and computationally analyzed. The work is concluded in Section 7, and some current research directions are outlined.
@@ 2 Notation, Abstract Setting, Finite Element Discretization
Throughout this work, let $d$ be the space dimension. We denote by $Q:=\Omega\times I\subset\mathbb{R}^{d+1}$ the space-time domain (cylinder), where $\Omega\subset\mathbb{R}^{d}$ is the bounded and Lipschitz spatial domain with the boundary $\Gamma=\partial\Omega$, and $I:=(0,T)$ denotes the temporal domain (time interval). Here, $T>0$ is the end time value. Furthermore, we use the standard notations for Lebesgue, Sobolev, and Bochner spaces like $L_{p}(\Omega)$, $W_{p}^{k}(\Omega),\mathring{W}_{p}^{k}(\Omega),H^{k}(\Omega)=W_{2}^{k}(\Omega),H^{k}_{0}(\Omega)=\mathring{W}_{2}^{k}(\Omega)$, $L_{p}(0,T;\mathring{W}_{p}^{1}(\Omega))$ etc.; see, e.g., [50, 52, 222].
@@ 2.2.2 Finite Elements on Hypercubes $Q_{k}$
To facilitate adaptive mesh refinement and to avoid connecting elements, we use the concept of hanging nodes. Elements are allowed to have nodes that lie on the midpoints of the faces or edges of neighboring cells. In our implementations, at most, one hanging node is allowed on each face or edge. In three dimensions, this concept is generalized to subplanes and faces because we must deal with two types of lower manifolds. To enforce global continuity (i.e., global conformity), the degrees of freedom located on the interface between different refinement levels have to satisfy additional constraints. They are determined by interpolation of neighboring degrees of freedom. Therefore, hanging nodes do not carry any degrees of freedom. For more details on this, we refer to [57].
@@ 3.2 Adjoint Problem
In order to realize goal-oriented error control in this work, we use the DWR method, which requires solving an adjoint problem. This adjoint formulation follows from the Lagrange formalism and is given by: Find $z\in V$ such that
@@ 3.3 An Error Identity
$\rho^{*}(\tilde{u},\tilde{z})(\cdot)$ $:=J^{\prime}(\tilde{u})-\mathcal{A}^{\prime}(\tilde{u})(\cdot,\tilde{z}),$ (7)
@@
$\mathcal{L}^{\prime}(\tilde{x})(x-\tilde{x})=$ $\underbrace{J^{\prime}(\tilde{u})(e)-\mathcal{A}^{\prime}(\tilde{u})(e,\tilde{z})}_{=\rho^{*}(\tilde{u},\tilde{z})(u-\tilde{u})}\underbrace{-A(\tilde{u})(e^{*})}_{=\rho(\tilde{u})(z-\tilde{z})},$
@@ 3.4 Error Estimation in Enriched Spaces
$\begin{split}\mathcal{R}^{(3)(2)}:=\frac{1}{2}\int_{0}^{1}&[J^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)})-\mathcal{A}^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)},\tilde{z}+se_{h}^{*(2)})\\ &-3\mathcal{A}^{\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{*(2)})]s(s-1)\,ds,\end{split}$ (13)
@@
holds, where $\rho(\tilde{u})(\cdot):=-\mathcal{A}(\tilde{u})(\cdot)$ and $\rho^{*}(\tilde{u},\tilde{z})(\cdot):=J^{\prime}(\tilde{u})-\mathcal{A}^{\prime}(\tilde{u})(\cdot,\tilde{z})$.
@@
where $\rho_{h}(\tilde{u})(\cdot):=-\mathcal{A}^{(2)}(\tilde{u})(\cdot)$, $\rho^{*}_{h}(\tilde{u},\tilde{z})(\cdot):=J^{\prime}_{h}(\tilde{u})-(\mathcal{A}^{(2)})^{\prime}(\tilde{u})(\cdot,\tilde{z})$, and
$\begin{split}\mathcal{R}_{h}^{(3)}:=\frac{1}{2}\int_{0}^{1}&[J_{h}^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)})-(\mathcal{A}^{(2)})^{\prime\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{(2)},\tilde{z}+se_{h}^{*(2)})\\ &-3(\mathcal{A}^{(2)})^{\prime\prime}(\tilde{u}+se_{h}^{(2)})(e_{h}^{(2)},e_{h}^{(2)},e_{h}^{*(2)})]s(s-1)\,ds,\end{split}$
@@ 3.5.3 The Discretization Error Estimator $\eta_{h}^{(2)}$
$\frac{1}{1+b_{h}}\left(|\eta_{h}^{(2)}|-|\eta_{k}+\eta_{\mathcal{R}}^{(2)}|\right)\leq|J(u)-J(\tilde{u})|\leq\frac{1}{1-b_{h}}\left(|\eta_{h}^{(2)}|+|\eta_{k}+\eta_{\mathcal{R}}^{(2)}|\right)$
@@
$\leq\frac{1}{1-b_{h}}|\eta^{(2)}|\leq\frac{1}{1-b_{h}}\left(|\eta_{h}^{(2)}|+|\eta_{k}+\eta_{\mathcal{R}}^{(2)}|\right).$
@@
$(1-\alpha_{\eta_{h}^{(2)}})\frac{1}{1+b_{h}}|\eta_{h}^{(2)}|\leq|J(u)-J(\tilde{u})|\leq(1+\alpha_{\eta_{h}^{(2)}})\frac{1}{1-b_{h}}|\eta_{h}^{(2)}|,$
@@ 3.6 Effectivity Indices for Enriched Spaces
$\frac{1}{1+b_{h}}|\eta^{(2)}|\leq|J(u)-J(\tilde{u})|\leq\frac{1}{1-b_{h}}|\eta^{(2)}|.$
@@
$\frac{1}{1+b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|\leq 1\leq\frac{1}{1-b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|.$
@@
$(1-\alpha_{\eta_{h}^{(2)}})\frac{1}{1+b_{h}}|\eta_{h}^{(2)}|\leq|J(u)-J(\tilde{u})|\leq(1+\alpha_{\eta_{h}^{(2)}})\frac{1}{1-b_{h}}|\eta_{h}^{(2)}|,$
@@
$\frac{1-\alpha_{\eta_{h}^{(2)}}}{1+b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|\leq 1\leq\frac{1+\alpha_{\eta_{h}^{(2)}}}{1-b_{h}}\left|\frac{\eta^{(2)}}{J(u)-J(\tilde{u})}\right|.$